# Bottom-up Policy Optimization: Your Language Model Policy Secretly Contains Internal Policies

### v1

@@ Abstract
Existing reinforcement learning (RL) approaches treat large language models (LLMs) as a single unified policy, overlooking their internal mechanisms. Understanding how policy evolves across layers and modules is therefore crucial for enabling more targeted optimization and raveling out complex reasoning mechanisms. In this paper, we decompose the language model policy by leveraging the intrinsic split of the Transformer residual stream and the equivalence between the composition of hidden states with the unembedding matrix and the resulting samplable policy. This decomposition reveals Internal Layer Policies, corresponding to contributions from individual layers, and Internal Modular Policies, which align with the self-attention and feed-forward network (FFN) components within each layer. By analyzing the entropy of internal policy, we find that: (a) Early layers keep high entropy for exploration, top layers converge to near-zero entropy for refinement, with convergence patterns varying across model series. (b) LLama’s prediction space rapidly converges in the final layer, whereas Qwen-series models, especially Qwen3, exhibit a more human-like, progressively structured reasoning pattern. Motivated by these findings, we propose Bottom-up Policy Optimization (BuPO), a novel RL paradigm that directly optimizes the internal layer policy during early training. By aligning training objective at lower layer, BuPO reconstructs foundational reasoning capabilities and achieves superior performance. Extensive experiments on complex reasoning benchmarks demonstrates the effectiveness of our method.
@@
[…]
Figure 1: (a): The residual stream within Transformer which moves from previous layer hidden states into self-attention and feed-forward network (FFN) sequentially. (b): Any hidden states with unembedding matrix $\mathbf{E}_{\text{u}}$ can be transformed into probability distribution $\mathbf{P}$ over the vocabulary space, which can be considered as the samplable policy. (c): We surprisingly find that Qwen-series contains a progressive reasoning pattern in FFN, where start from exploration expansion to integrate middle layer knowledge into final prediction convergence, specially in Qwen3.
@@ 1 Introduction
Reinforcement learning (RL) has emerged as a key driver in advancing the complex reasoning capabilities of large language models (LLMs). Notably, the success of DeepSeek R1 (guo2025deepseek) has solidified reinforcement learning with verifiable rewards (RLVR) as a potent post-training paradigm for enhancing LLM policies across diverse domains (jaech2024openai; yang2025qwen3; team2025kimi; ouyang2022training). To date, most existing RLVR research has predominantly focused on surface algorithmic design, such as reward construction (shao2025spurious; chen2025pass; chen2025seed; liu2025understanding) and entropy regularization (cui2025entropy; yu2025dapo; yang2025dcpo).
Interpretability tools mitigate the opacity of black-box LLMs by unveiling their internal logic (belrose2023eliciting; tan-etal-2025-neural; gupta2025llms). These insights establish a feedback loop that is instrumental in guiding the refinement of RL algorithms. For instance, recent works leverage the interpretability tool of attention mechanisms to improve RL algorithm which show surprising effect (li2025attention; liu2025attention). However, they overlook the nature of the language model policy as well as the information latent in the model’s internal residual stream.
@@
The logit lens (nostalgebraist2020logitlens) framework offers initial insights by employing the unembedding matrix $\mathbf{E}_{\text{u}}$ to decode intermediate layer representations into the token space. This approach reveals that the internal residual stream harbors a wealth of previously undiscovered information that evolves across layers and modules (dai2022knowledge; gupta2025llms; lindsey2025biology). Moreover, numerous studies have elucidated the mechanisms of self-attention and feed-forward networks (FFNs) and their impact on hidden states (dai2022knowledge; yu2023neuron; jin2025massive). Collectively, these internal mechanisms with models offer a new perspective for algorithmic optimizations.
@@
In this paper, we investigate the evolution of language model policies across layers and modules to facilitate optimization and unravel complex internal reasoning mechanisms. Our formulation is grounded in two key insights. First, the residual stream naturally supports additive decomposition (zhang2025reinforcement; lindsey2025biology), allowing us to isolate the individual roles of each layer and module (Figure 1(a)). Second, we conclude that the samplable policy is intrinsic equivalent to the token distribution derived from the combination of hidden states $\mathbf{H}$ with unembedding matrix $\mathbf{E}_{\text{u}}$. Based on these, we construct the Internal Layer Policy $\pi_{\text{Layer}}^{l}$, which captures cumulative reasoning up to layer $l$, and the Internal Modular Policy $\pi_{\text{ATTN}}^{l}$ and $\pi_{\text{FFN}}^{l}$, which isolates the specific contributions of attention and FFN modules (Figure1(b)). This decomposition allows us to ask: How does internal reasoning evolve through the model?
@@
Through systematic analysis of commonly used Qwen and Llama series (meta2024llama32; yang2024qwen2; yang2025qwen3) based on Internal Policy Entropy $H$ in a policy-centric view, we uncover both universal and critical architectural differences: (1) Consistent internal reasoning structure. All models exhibit a universal reasoning structure: early layers maintain high entropy for exploring the solution space, while top layers converge to near-zero entropy for final prediction (lindsey2025biology). (2) Distinct internal reasoning pattern. Despite the shared trend, the pace of convergence differs significantly. Llama exhibits a sudden convergence, where the prediction space collapses only within the last three layers. In contrast, Qwen models demonstrate a progressive contraction, gradually reducing uncertainty throughout the layers. To quantify these dynamics, we introduce Internal Entropy Change $\Delta H$. This metric reveals that while Llama shows minimal intermediate updates, Qwen (especially Qwen3) utilizes FFNs in a progressive, human-like cognitive process (dehaene1998neuronal): expanding exploration in lower layers, integrating parametric knowledge in intermediate layers (dai2022knowledge), and compacting predictions in upper layers (Figure 1(c)).
@@
These findings have profound implications for RL optimization: Since internal reasoning emerges progressively from lower to higher, we can consider optimization from a bottom-up perspective. We initially validate this hypothesis through targeted optimization experiments on internal policies, revealing distinct training dynamics and a remarkable phenomenon of internal reasoning feature refinement. Specifically, the optimized lower layers capture high-level reasoning capabilities by early alignment, providing a more robust foundation for subsequent internal reasoning. Motivated by these insights, we propose Bottom-up Policy Optimization (BuPO), a novel reinforcement learning paradigm that optimizes fine-grained internal layer policies during the early stages of training to effectively guide the overall language model policy. By doing so, BuPO reconstructs foundational reasoning abilities and achieves superior performance. Extensive experiments on complex reasoning benchmarks demonstrate the effectiveness of our approach and the unique training dynamics compared to conventional RL methods that optimize the policy as a whole. We hope our detailed analysis of language model policies provides valuable insights to improve RL algorithms.
In summary, this paper makes the following contributions:
  • Internal Policy Decomposition: First, we formally define and decompose language model policies into internal layer and modular policies, revealing their distinct roles in reasoning.
  • Internal Policy Entropy-based Analysis: Building on this foundation, we uncover consistent yet distinct reasoning patterns across different model series, highlighting progressive reasoning structures in Qwen models.
  • Bottom-up Policy Optimization: Inspired by these findings, we propose BuPO, a novel RL paradigm that optimizes internal policies early in training, achieving superior performance and unique training behaviors.
@@ 2 Preliminary
2.1 The Information Flow in Transformer
Transformer-based language models (vaswani2017attention) form the foundation of modern LLMs (brown2020language). A decoder-only Transformer consists of $L$ stacked layers, each containing a multi-head self-attention (MHSA) module and a feed-forward network (FFN) module.
Following yu2023neuron; zhang2025reinforcement, we formalize the forward process from input to output. Given an input sequence $\mathbf{x}=[x_{1},x_{2},\dots,x_{T}]$, the model produces a probability distribution $\mathbf{P}$ over the vocabulary $V$ with $N$ tokens. Let $\mathbf{H}^{(2l-2)}\in\mathbb{R}^{T\times d_{\text{model}}}$ denote the hidden state input to the $l$-th layer, where $S$ is the sequence length and $d_{\text{model}}$ is the hidden dimension. The initial embedding is $\mathbf{H}^{(0)}$ projected by $\mathbf{E}$, where $\mathbf{E}\in\mathbb{R}^{N\times d_{\text{model}}}$ is the embedding matrix.
Each layer processes information sequentially through attention and FFN:
@@ 2.2 Reinforcement Learning for Language Model Policy
(a) Continuous entropy dynamics of internal policy for different models. The residual stream flows from $\mathbf{H}^{l-1}$ into $\mathbf{A}^{l}$, $\mathbf{F}^{l}$, and finally to the next layer $\mathbf{H}^{l}$.
@@
where $R(\mathbf{q},\mathbf{o})=\sum_{t=1}^{|\mathbf{o}|}r(s_{t},o_{t})$ is the return (sutton1998reinforcement) and $\pi_{\text{ref}}$ is a reference policy. We adopt sparse rewards where $r_{t}=0$ for $t<n$ and $r_{n}\in[0,1]$ indicates task success. Following hu2025openreasonerzeroopensourceapproach, we will assume $\beta=0$ throughout this paper.
Policy Optimization. We adopt GRPO (shao2024deepseekmath), which samples a group of responses $\{\mathbf{o}_{1},\dots,\mathbf{o}_{G}\}$ per question and estimates advantages as $\hat{A}_{i,t}=\frac{R_{i}-\operatorname{mean}(\mathbf{R})}{\operatorname{std}(\mathbf{R})}$:
@@ 3.1 Definition of Internal Policy
Residual stream. Due to residual connections in Transformer layers, the input to any sub-module (attention or FFN) equals the sum of all preceding outputs plus the original embedding. Let $\mathbf{A}^{l}=\mathcal{A}^{l}(\mathbf{X}^{l}_{\text{attn}})$ and $\mathbf{F}^{l}=\mathcal{F}^{l}(\mathbf{X}^{l}_{\text{ffn}})$ denote attention and FFN transformations (excluding normalization). The hidden states satisfy:
@@
Internal policy. During reinforcement tuning, we sample the next token $o_{t}$ from the final layer’s probability distribution, i.e., $\pi_{\theta}\equiv\mathbf{P}=\text{softmax}(\texttt{LN}(\mathbf{H}^{L})\mathbf{E}_{\text{u}}^{\text{T}})$. We propose that every internal hidden state can combine with $\mathbf{E}_{\text{u}}$ to produce a samplable internal policy. Specifically, we focus on two granularities. Internal Layer Policy refers to utilize hidden states from each layer $\mathbf{H}^{l}$ to combine with $\mathbf{E}_{\text{u}}$, and Internal Modular Policy integrates $\mathbf{E}_{u}$ with hidden states from specific module:
@@
Each component contributes to the final policy through the residual stream. For instance, $\mathbf{H}^{L}=\mathbf{H}^{l}+\mathbf{S}^{l+1}$, where $\mathbf{S}^{l+1}=\sum^{L}_{i=l+1}\mathbf{A}^{i}+\sum^{L}_{j=l+1}\mathbf{F}^{j}$ represents contributions from subsequent layers. Hence, understanding these internal components is essential for raveling how does reasoning emerge and evolve through the model.
@@ 3.2 Internal Policy Entropy Dynamics
In contrast to prior logit lens approaches (nostalgebraist2020logitlens; belrose2023eliciting) that decode internal states into discrete tokens, we adopt a policy-centric perspective where internal probability distributions can be sampled and updated like policies. We employ entropy as our primary metric, motivated by its strong correlation with policy behavior in RL (cui2025entropy; agarwal2025unreasonable; cheng2025reasoning). We define Internal Policy Entropy as:
@@
Continuous entropy dynamics. Figure 7(a) shows that internal policy entropy dynamics exhibit consistent patterns across models: early layers maintain high entropy for exploration of the search space, while top layers converge to near-zero entropy. This aligns with findings that lower layers capture semantic information while higher layers aggregate and refine these representations to drive final decision-making (lindsey2025biology).
While the overall entropy pattern is consistent across models, the fine-grained module-wise transition dynamics vary. To isolate intrinsic patterns from normalization and residual effects (zhang2019root; he2016deep)), we introduce Entropy Change, which measures the incremental information gain within a single internal policy and is defined as:
$\Delta H^{l}=H^{l}_{\text{Output}}-H^{l}_{\text{Input}}$ (9)
where the entropy change is defined as the difference between the internal policy entropy at the input and output of a given module, e.g., $\Delta H^{l}_{\text{FFN}}$. This quantity characterizes how the exploration space evolves as information propagates through the module. Specifically, $\Delta H^{l}_{\text{FFN}}>0$ indicates an expansion of exploration, $\Delta H^{l}_{\text{FFN}}\approx 0$ suggests internal knowledge integration, and $\Delta H^{l}_{\text{FFN}}<0$ reflects prediction convergence during the reasoning process.
Entropy change dynamics of attention vs. FFN. Self-attention modules (vaswani2017attention) are widely regarded as central to model reasoning, particularly for integrating task-relevant contextual information (jin2025massive; liu2025attention). The upper panel of Figure 7 reveals a clear and model-dependent pattern in the entropy change of self-attention. Specifically, Qwen3 models exhibit consistently positive entropy change across layers ($\Delta H^{l}_{\text{ATTN}}>0$), indicating sustained expansion of the exploration space during reasoning, consistent with prior findings (li2025attention; lindsey2025biology; zhou2024role). In contrast, Qwen2.5-Math-7B shows uniformly negative entropy change, suggesting progressive contraction and earlier convergence. Llama models display a weaker but still positive trend, reflecting more conservative exploration dynamics.
In contrast, the FFN module is widely regarded as the key-value memories of parametric knowledge in Transformers (geva2021transformer; meng2022locating; dai2022knowledge). The lower panel of Figure 7 reveals clear and systematic differences in FFN entropy dynamics across model families. For the Llama-series models (meta2024llama32), FFN entropy remains consistently positive across almost all layers, with convergence occurring only at the final layer. This pattern indicates sustained exploration throughout the FFN, with limited intermediate consolidation, which aligns with prior observations that Llama models benefit less from post-training and require additional mid-training interventions to improve reasoning behavior (wang2025octothinker).
By contrast, the Qwen-series models (yang2024qwen2; yang2025qwen3) exhibit a pronounced hierarchical entropy structure in the FFN, following a clear three-stage progression. Taking Qwen3-4B as an illustrative example, the lower FFN layers (layers 1–6) show increasing entropy ($\Delta H^{l}_{\text{FFN}}>0$), corresponding to expanded exploration at the onset of reasoning. This is followed by a broad middle region (layers 7–26) where entropy change remains approximately zero, indicating stable information integration through retrieval and reuse of parametric knowledge encoded in knowledge neurons (dai2022knowledge). In the upper layers (layers 27–36), entropy decreases ($\Delta H^{l}_{\text{FFN}}<0$), reflecting gradual convergence toward the final prediction. Figure 1(c) provides a vivid demonstration.
Notably, this “Exploration–Integration–Convergence” (EIC) pattern is more consistently expressed in Qwen3 than in earlier Qwen2.5 variants, indicating a more structurally stabilized and progressive reasoning pattern. Rather than maintaining continuous exploration across all layers, Qwen3 allocates distinct FFN depth ranges to different reasoning functions, resulting in a staged reasoning process. We hypothesize that this progressive reasoning pattern, qualitatively similar to the staged nature of human reasoning (dehaene1998neuronal; felleman1991distributed; yamins2016using), may help explain why Qwen3 exhibits more efficient knowledge absorption during post-training (zhu2025surprising; yang2025qwen3).
(a) Residual cosine similarity across different Qwen models

### v2

@@ Abstract
Existing reinforcement learning (RL) approaches treat large language models (LLMs) as a unified policy, overlooking their internal mechanisms. In this paper, we decompose the LLM-based policy into Internal Layer Policies and Internal Modular Policies via Transformer’s residual stream. Our entropy analysis on internal policy reveals distinct patterns: (1) universally, policies evolve from high-entropy exploration in early layers to deterministic refinement in top layers; and (2) Qwen exhibits a progressive, human-like reasoning structure, contrasting with the abrupt final-layer convergence in Llama. Furthermore, we discover that optimizing internal layers induces feature refinement, forcing lower layers to capture high-level reasoning representations early. Motivated by these findings, we propose Bottom-up Policy Optimization (BuPO), a novel RL paradigm that reconstructs the LLM’s reasoning foundation from the bottom up by optimizing internal layers in early stages. Extensive experiments on complex reasoning benchmarks demonstrate the effectiveness of BuPO.
@@
tanyuqiao2025@ia.ac.cn,wangminzheng2023@ia.ac.cn, shizhu.he@nlpr.ia.ac.cn
Figure 1: (a): The residual stream within Transformer moves from previous layer into self-attention and feed-forward network (FFN) sequentially. (b): Any hidden states with unembedding matrix $\mathbf{E}_{\text{u}}$ can be transformed into probability distribution $\mathbf{P}$ over the vocabulary space, which can be considered as a policy. (c): We surprisingly find that Qwen-series contains a progressive reasoning pattern in FFN, where it starts from exploration expansion to integrate middle layer knowledge into final prediction convergence.
@@ 1 Introduction
Reinforcement learning (RL) has emerged as a key driver in advancing the complex reasoning capabilities of large language models (LLMs) (Ouyang et al., 2022; Jaech et al., 2024). Notably, the success of DeepSeek-R1 (Guo et al., 2025) has solidified reinforcement learning with verifiable rewards (RLVR) as a potent post-training paradigm for enhancing language model policy across diverse domains (Yang et al., 2025; Team et al., 2025; Yu et al., 2025). However, existing RL approaches typically treat the LLM as a unified policy, optimizing it solely on the final output distribution while neglecting the intricate information flow evolving through its internal residual streams. Mechanistic interpretability tools mitigate the opacity of black-box LLMs by unveiling their internal logic (Belrose et al., 2023; Tan et al., 2025a; Gupta et al., 2025). Such insights are viewed as a crucial bridge between deeper understanding of internal behavior and principled algorithmic refinement.
@@
While recent works leverage the attention mechanisms to improve RL algorithm (Li et al., 2025; Liu et al., 2025), they overlook the nature of the language model policy as well as the information latent in the residual stream. The logit lens (nostalgebraist, 2020) framework offers initial insights by employing the unembedding matrix $\mathbf{E}_{\text{u}}$ to decode intermediate layer representations into the token space. This approach reveals that the internal residual stream harbors a wealth of undiscovered information that evolves across layers and modules (Dai et al., 2022; Gupta et al., 2025; Lindsey et al., 2025). Moreover, numerous studies have elucidated the mechanisms of self-attention and feed-forward network (FFN)  (Dai et al., 2022; Yu and Ananiadou, 2024; Jin et al., 2025). Collectively, these mechanisms within models offer a new perspective for algorithmic optimizations.
@@
In this paper, we investigate the evolution of language model policies across layers and modules to facilitate optimization and unravel complex internal reasoning mechanisms. Our formulation is grounded in two key insights. First, the residual stream naturally supports additive decomposition (Zhang et al., 2025; Lindsey et al., 2025), allowing us to isolate the individual roles of each layer and module (Figure 1(a)). Second, we conclude that the samplable policy is intrinsic equivalent to the token distribution derived from the combination of hidden states $\mathbf{H}$ with unembedding matrix $\mathbf{E}_{\text{u}}$. Based on these, we construct the Internal Layer Policy $\pi_{\text{Layer}}^{l}$, which captures cumulative reasoning up to layer $l$, and the Internal Modular Policy $\pi_{\text{ATTN}}^{l}$ and $\pi_{\text{FFN}}^{l}$, which isolates the specific contributions of attention and FFN modules (Figure 1(b)). This decomposition allows us to ask: How does internal reasoning evolve through the model?
@@
Through systematic analysis of commonly used Qwen and Llama series (Meta AI, 2024; Yang et al., 2024, 2025) based on Internal Policy Entropy in a policy-centric view, we uncover both universal and critical architectural differences: (1) Consistent internal reasoning structure. All models exhibit a universal reasoning structure: early layers maintain high entropy for exploring the solution space, while top layers converge to near-zero for final prediction (Lindsey et al., 2025). (2) Distinct internal reasoning pattern. Despite the shared trend, the pace of convergence differs significantly. Llama exhibits a sudden convergence only within the last three layers. In contrast, Qwen models demonstrate a progressive contraction, gradually reducing uncertainty throughout layers. To quantify these dynamics, we introduce Entropy Change. This metric reveals that while Llama shows minimal internal updates, Qwen utilizes FFN in a progressive, human-like cognitive process (Dehaene et al., 1998): expanding exploration in lower layers, integrating parametric knowledge in intermediate layers, and compacting predictions in upper layers (Figure 1(c)).
@@
These findings have profound implications for RL optimization: Since internal reasoning emerges progressively from lower to higher, we can consider optimization from a bottom-up perspective. We initially validate this hypothesis through targeted optimization on internal policies, revealing distinct training dynamics and a remarkable phenomenon of internal reasoning feature refinement. Specifically, the optimized lower layers capture high-level reasoning capabilities by early alignment, providing a more robust foundation for subsequent internal reasoning. Motivated by these insights, we propose Bottom-up Policy Optimization (BuPO), a novel RL paradigm that optimizes fine-grained internal layer policies during the early stages of training to effectively guide the overall language model policy. By doing so, BuPO reconstructs foundational reasoning abilities and achieves superior performance. Extensive experiments on complex reasoning benchmarks demonstrate the effectiveness of our approach and the unique training dynamics compared to conventional RL methods that optimize the policy as a whole.
In summary, this paper makes the following contributions: (1) We are the first to decompose an LLM policy into internal layer and modular policies , revealing their distinct roles in reasoning. (2) We reveal a universal exploration-to-convergence shift, distinguishing Qwen’s progressive reasoning from Llama’s abrupt convergence. Moreover, we uncover a feature refinement phenomenon where internal optimization drives lower layers to preemptively capture high-level features. (3) We propose BuPO to align internal policies from the bottom up, which reconstructs the foundational reasoning capabilities at lower layers, achieving superior performance on complex reasoning benchmarks.
(a) Continuous entropy dynamics of internal policy. The residual stream flows from $\mathbf{H}^{l-1}$ into $\mathbf{A}^{l}$, $\mathbf{F}^{l}$, then to the next layer $\mathbf{H}^{l}$.
@@ 2 Preliminary
2.1 The Residual Stream in Transformer
Transformer-based language models (Vaswani et al., 2017) form the foundation of modern LLMs (Brown et al., 2020). A decoder-only Transformer consists of $L$ stacked layers, each containing a multi-head self-attention (MHSA) module and a feed-forward network (FFN) module.
Following Zhang et al. (2025), we formalize the forward process from input to output. Given an input sequence $\mathbf{x}=[x_{1},x_{2},\dots,x_{T}]$, the model produces a probability distribution $\mathbf{P}$ over the vocabulary $V$ with $N$ tokens. Let $\mathbf{H}^{(2l-2)}\in\mathbb{R}^{T\times d_{\text{model}}}$ denote the hidden state input to the $l$-th layer, where $T$ is the sequence length and $d_{\text{model}}$ is the hidden dimension. The initial embedding is $\mathbf{H}^{(0)}$ projected by $\mathbf{E}$, where $\mathbf{E}\in\mathbb{R}^{N\times d_{\text{model}}}$ is the embedding matrix.
Each layer forwards sequentially through attention and FFN:
@@ 2.2 Reinforcement Learning for Language Model Policy
where $R(\mathbf{q},\mathbf{o})=\sum_{t=1}^{|\mathbf{o}|}r(s_{t},o_{t})$ is the return (Sutton et al., 1998) and $\pi_{\text{ref}}$ is a reference policy. We adopt sparse rewards where $r_{t}=0$ for $t<n$ and $r_{n}\in[0,1]$ indicates task success. Following Hu et al. (2025a), we assume $\beta=0$.
Policy Optimization.
We adopt GRPO (Shao et al., 2024), which samples a group of responses $\{\mathbf{o}_{1},\dots,\mathbf{o}_{G}\}$ per question and estimates advantages as $\hat{A}_{i,t}=\frac{R_{i}-\operatorname{mean}(\mathbf{R})}{\operatorname{std}(\mathbf{R})}$:
@@ 3.1 Definition of Internal Policy
Residual Stream.
In the residual stream of Transformer, the input to any layer equals the sum of all preceding outputs plus the original embedding. The hidden states satisfy:
@@
Internal Policy.
During RL, we sample the next token $o_{t}$ from the final layer’s probability distribution, i.e., $\pi_{\theta}\equiv\mathbf{P}=\text{softmax}(\texttt{LN}(\mathbf{H}^{L})\mathbf{E}_{\text{u}}^{\text{T}})$. We propose that every internal hidden states can combine with $\mathbf{E}_{\text{u}}$ to produce a samplable policy. Specifically, we focus on two granularities. Internal Layer Policy refers to utilize hidden states from each layer $\mathbf{H}^{l}$ to combine with $\mathbf{E}_{\text{u}}$, and Internal Modular Policy integrates $\mathbf{E}_{u}$ with hidden states from specific module:
@@
Each component contributes to the final policy through the residual stream. For instance, $\mathbf{H}^{L}=\mathbf{H}^{l}+\mathbf{S}^{l+1}$, where $\mathbf{S}^{l+1}=\sum^{L}_{i=l+1}\mathbf{A}^{i}+\sum^{L}_{j=l+1}\mathbf{F}^{j}$ represents contributions from subsequent layers. Hence, understanding these internal components is essential for unraveling how internal reasoning emerge and evolve through the model.
@@ 3.2 Internal Policy Entropy Dynamics
In contrast to prior logit lens approaches (nostalgebraist, 2020; Belrose et al., 2023) that decode internal states into discrete tokens, we adopt a policy-centric perspective where internal probability distributions are treated as policies. We employ Entropy as our primary metric, motivated by its strong correlation with policy behavior (Cui et al., 2025; Cheng et al., 2025). We define Internal Policy Entropy as:
@@
Continuous Entropy Dynamics.
Figure 7(a) shows that internal policy entropy dynamics exhibit consistent patterns across models: early layers maintain high entropy for exploration of the search space, while top layers converge to near-zero entropy. This aligns with findings that lower layers capture semantic information while higher layers aggregate and refine these representations to drive final decision-making (Lindsey et al., 2025).
While the overall entropy pattern is consistent across models, the fine-grained transition dynamics vary. To isolate intrinsic patterns from normalization and residual effects (He et al., 2016; Zhang and Sennrich, 2019), we introduce Entropy Change, which measures the incremental information gain within a single internal policy and is defined as:
$\Delta H^{l}=H^{l}_{\text{Output}}-H^{l}_{\text{Input}},$ (9)
where the entropy change is defined as the difference between the internal policy entropy at a module’s input and output. This metric reveals how the exploration space evolves as information propagates through the specific module.
@@
Entropy Change Dynamics of Attention vs. FFN.
Self-attention modules (Vaswani et al., 2017) are widely regarded as central to model reasoning, particularly for integrating task-relevant contextual information (Jin et al., 2025; Liu et al., 2025). The upper panel of Figure 7 reveals a clear and model-dependent pattern in the entropy change of self-attention. Specifically, Qwen3 models exhibit consistently positive entropy change across layers ($\Delta H^{l}_{\text{ATTN}}>0$), indicating sustained expansion of the exploration space during reasoning, consistent with prior findings (Li et al., 2025; Lindsey et al., 2025; Zhou et al., 2025). In contrast, Qwen2.5-Math-7B shows uniformly negative entropy change, suggesting progressive contraction and earlier convergence. Llama models display a weaker but still positive trend, reflecting more conservative exploration dynamics.
Moreover, the FFN module is widely regarded as the key-value memories of parametric knowledge (Geva et al., 2021; Meng et al., 2022; Dai et al., 2022). The lower panel of Figure 7 reveals clear and systematic differences in FFN entropy dynamics across model families. For the Llama models (Meta AI, 2024), FFN entropy remains consistently positive across almost all layers, with convergence occurring only at the final layer. This pattern indicates sustained but weak exploration throughout the FFN, which aligns with prior observations that Llama models benefit less from post-training and require additional mid-training interventions to improve reasoning behavior (Wang et al., 2025b).
By contrast, the Qwen models (Yang et al., 2024, 2025) exhibit a pronounced hierarchical entropy structure in the FFN, following a clear three-stage progression. Taking Qwen3-4B as an example, the lower FFN layers (layers 1–6) show increased entropy ($\Delta H^{l}_{\text{FFN}}>0$), corresponding to expanded exploration at the onset of reasoning. Followed by a broad middle region (layers 7–26) where $\Delta H^{l}_{\text{FFN}}\approx 0$, indicating stable information integration through retrieval and reuse of parametric knowledge  (Dai et al., 2022). In the upper layers (layers 27–36), $\Delta H^{l}_{\text{FFN}}<0$, reflecting gradual convergence toward the final prediction.
Notably, this Exploration–Integration–Convergence (EIC) pattern is more consistently expressed in Qwen3 than in earlier Qwen2.5 variants, indicating a more structurally stabilized and progressive reasoning pattern. We hypothesize that this progressive reasoning pattern, qualitatively similar to the staged nature of human reasoning (Dehaene et al., 1998; Yamins and DiCarlo, 2016), may help explain why Qwen3 exhibits more efficient knowledge absorption during post-training (Zhu et al., 2025; Yang et al., 2025). The analysis of residual stream similarity in Appendix B.2 further supports our finding in EIC pattern.
4 Internal Policy Optimization
Building on previous analysis, we observe that internal reasoning emerges progressively, with different internal policy regions serving distinct roles and exhibiting varied behaviors. This raises a key question: Can we approach bottom-up optimization to align with this progressive emergence? To address this, we extend GRPO which designed for language model policy to directly optimize these internal policies:
$\frac{1}{G}\sum^{G}_{i=1}\frac{1}{|\mathbf{o}_{i}|}\sum_{t=1}^{|\mathbf{o_{i}}|}\left\{\min\left[\hat{r}_{i,t}\hat{A}_{i,t},\text{clip}(\hat{r}_{i,t},1-\epsilon,1+\epsilon)\hat{A}_{i,t}\right]\right\},$
where we sample from $\pi_{\theta_{\text{old}}}$ and $\hat{r}_{i,t}=\frac{\pi_{\text{Layer}}^{l}(o_{i,t}|\mathbf{q},\mathbf{o}_{i,<t})}{\pi^{l}_{\text{Layer,old}}(o_{i,t}|\mathbf{q},\mathbf{o}_{i,<t})}$ for current optimizing policy $\pi^{l}_{\text{Layer}}$ and keep other designs same with Eq. 4. We use Qwen3-4B as the backbone and present the implementation details in Appendix A.5.